calculate the no. of moles and the amount in grams of NaOH in 250 cubic cm. of a 0.100 M NaOH solution?

Grade:

2 Answers

Rahul askIITiansExpert.IITR
13 years ago

0.1 M solution has 0.1 moles of NaOH in 1 L of solution that is to say 1000 cubic cm.

so 250 cubic cm. solution will have 0.1/4 that is to say .025 moles of NaOH.

Now multiply the number of moles by the Molecular weight of NaOH to get the amoun in grams.

Please feel free to post as many doubts on our discussion forum as you can. If you find any question

Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation.

DO APPROVE IF THE ANSWER HELPS YOU.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Regards,

Askiitians Experts

Rahul- IIT Roorkee

ujjwal kumar
13 years ago

very simple ques... but ok. here is ur sol.

1 cubic cm = 1/1000 litre..... so 250 cubic cm = 250/1000,  "means 0.25 liter. NaOH present"
given sol. of 0.100 M.  therefore total mole present ; 0.100M*0.25= 0.025 ( 1st answer)

2nd answer...  1 mole contain 32 gram NaOH ( 23+8+1 )....  therefore amt of gram of NaOH present ; 0.025*32= calculate urself